OpenStudy (anonymous):

Anna's mother is three times as old as Anna. Ten years from now, she will only be twice as old as Anna. How old are they?

OpenStudy (ajprincess):

Let the present age of Anna be x. then her mother's age will be 3x. After 10 yrs Anna's age=x+10 Mother's age=3x+10 2(x+10)=3x+10

OpenStudy (ajprincess):

2(x+10)=3x+10 2x+20=3x+10 solve for x. then find 3x, Does that help? @jillang07

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

welcome:)